<p>Split funnels were given to Oemer for cleaning. Have received them back after lunch.</p>
<p>Valve support is installed again (see picture 1).</p>
<p>Ion pump was connected but not activated. Michael will help in doing so. No vacuum support to help will RGA. Will proceed without it.</p>
<p>Neg and ion cables were labeled (see pictures 2 and 3).</p>
<p>Chamber was vented using nitrogen.</p>
<p> </p> |